Solve the following system of equations using the substitution method
5x-7y= -18
4x+ 3y = 20

Answer by Cromlix(4381)

5x-7y= -18....(1)
4x+ 3y = 20...(2)
Altering (2)
4x + 3y = 20
4x = 20 - 3y
x = 5 - 3/4y
...........
Substituting in:-
5x - 7y = -18
5(5 - 3/4y) - 7y = -18
25 - 15/4y - 7y = -18
Convert 7 into 28/4
25 - 15/4y - 28/4y = - 18
25 - 43/4y = - 18
- 43/4y = -18 - 25
- 43/4y = - 43
43/4y = 43 (Multiply both sides by -1)
43y = 43 x 4
43y = 172
y = 172/43
y = 4
Substitute y = 4 into
4x + 3y = 20
4x + 3(4) = 20
4x + 12 = 20
4x = 20 - 12
4x = 8
x = 2
{2,4}
Hope this helps:-)